40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1690     Entry time: Thu Jul 28 19:56:08 2016
Author: Antonio, Andrew 
Type: DailyProgress 
Category: BEAT 
Subject: Searching for beat note 

Summary
In the last two days we have been searching for beat note, in order to get 
our "first" noise measurement of the PLL. While we monitor the beat note of the lasers
at the input side we search the same beat note at the transmission side.

Conclusions
We did not succeded yet. I start to believe that what we see at the input side are
not the beat note we are looking for. Some mode hopping could be happening around
the temperature region we have chosen.
 


Things done

1. We built a Mach Zehnder interferometer at the input side of the lasers, in order to fined the beat
directly from the two lasers; The beat note were found around 450MHz.

2. Once we found them we changed the temperature of the vacuum chamber and we locked the cavities
whit laser beat note of 78MHz.

3. Trasmitted beam have been checked for their spatial overlap and polarizations;

4. Photodiodes have been changed in order to see if some problem was coming from there;

5. Network analizer have been changed in order to check if there were wrong settings.

6. PLL rialigned from beginning;


NO RESULTS

PLAN 
We may want to know where mode hopping happen; We need to scan the temperature and  first measure
the power and see where they are. 

Looking for other beat note at the lasers.

ELOG V3.1.3-